OpenStudy (anonymous):

Calculate the frequency of radiation with a wavelength of 8973 (symbol following is A with a circle on top)

OpenStudy (whpalmer4):

\[8973{\buildrel _{\circ} \over {\mathrm{A}}} \]is that what the wavelength is? Do you know the definition of the \({\buildrel _{\circ} \over {\mathrm{A}}}\) unit?

OpenStudy (whpalmer4):

\[1\text{ } {\buildrel _{\circ} \over {\mathrm{A}}} = 10^{-10} \text{ m}\] Frequency and wavelength for electromagnetic waves are related by \[c = f \lambda\] where \(c\) is the speed of light in vacuum, \(2.99792458*10^8\text{ m/s}\) \(f\) is frequency and \(\lambda\) is wavelength in meters

OpenStudy (anonymous):

yes that is what I was trying to say, thank you I will get back to you if i have questions

OpenStudy (anonymous):

so the wavelength would be : 8973*10^-10?

OpenStudy (whpalmer4):

yes, wavelength is \(8973*10^{-10}\text{m}\) that gives you \[c = f \lambda\]\[\lambda = \frac{c}{f} = \frac{2.99792458*10^8\text{ m/s}}{8973*10^{-10}\text{m}} =\]

OpenStudy (anonymous):

0.000003341 would that be 3.3*10^14 Hz?

OpenStudy (whpalmer4):

I would carry 4 decimal places to match the 4 you had for the wavelength, but yes, \[3.341*10^{14} \text{ Hz}\]

OpenStudy (anonymous):

okay thank you so much, this helped a lot!!

OpenStudy (whpalmer4):

glad to hear it!